Kandi has $2.65 in nickels and dimes. If she has 40 coins in all, how many nickels does she have?
Found 2 solutions by AnlytcPhil, funmath:
Answer by AnlytcPhil(1806)


Kandi has $2.65 in nickels
and dimes. If she has 40 coins in all,
how many nickels does she have?


COIN equation:         N + D = 40

MONEY equation   .05N + .10D = 2.65

Can you solve that system of equations? 
If not post again.

Answer: N = 23, D = 13

Edwin

Answer by funmath(2933)

If N is nickles and D is dimes then N+D=40 because Kandi has 40 coins.
If we solve this equation for N we get:
N+D-N=40-N
D=40-N
So the number of dimes we have is 40-N
and the number of nickles we have is N
The amount of money we have is equal to the worth of the coin times the number of coins. I like to solve these equations in terms of cents rather than dollars because it eliminates the decimals.
The amount of money we have in dimes is 10(40-N).
The amount of money we have in nickles is 5N.
The total amount of money we have is 265 (in cents).
Therefore, (the amount of money in dimes)+(the amount of money in nickles)=(total) becomes:
10(40-N)+5N=265
400-10N+5N=265
400-5N=265
-400+400-5N=265-400
-5N=-135
-5N/-5=-135/-5
N=27
Thus, Kandi had 27 nickels.